# Scalable Data-Driven Basis Selection for Linear Machine Learning Interatomic Potentials

### v1

@@ I Introduction
For large linear MLIPs models, achieving a balance between model complexity, interpretability, and generalizability remains a challenge [23, 24]. Sparse selection of features can in principle lead to minimal yet informative descriptions. Such models tend to enhance interpretability, scalability, and generalization to unseen configurations. An early example of sparsification in this context appears in the MTP model [41], where a matching pursuit algorithm with crossover and local search was used to minimize an objective with an $\ell_{0}$ regularization term, to train a potential for tungsten. While sparsity was not the primary focus, this work demonstrated that a minimal basis could be identified without sacrificing accuracy. Automatic relevance determination (ARD) [33], an empirical Bayes method that can also be viewed as a sparsification technique, was used in the context of active learning of ACE models [50]. In the present work, we explore the use of two alternative sparse optimization methods, active set basis pursuit (ASP) [27], orthogonal matching pursuit (OMP) [27, 38], and compare them against ARD and various “dense” least squares solvers. The ASP and OMP methods we employ produce entire “paths” of MLIPs models with increasing model complexity. We apply those parameter estimation methods in the linear ACE framework [25].
Our computational tests on three benchmark problems demonstrate several advantages. First, the sparse solvers require minimal user intervention and substantially reduce the need for manual hyperparameter tuning. Secondly, they lead to improved generalization, demonstrated by improved test errors compared to non-sparse solvers. Finally, the selected features do not exhibit a clear or predictable pattern, emphasizing the benefit of data-driven feature selection over heuristic or pre-defined approaches.
II Methods
II-A Machine Learning Interatomic Potentials (MLIPs)
@@
A common approach to constructing interatomic potentials is to decompose the total energy $E$ of the system as a sum of the site energies: $E=\sum_{i}\epsilon_{i}.$ In linear models, each site energy $\epsilon_{i}$ is expressed as a linear combination of invariant basis functions that encode information about the local environment (e.g., interatomic distances, angles, etc.) within a cutoff radius $r_{\text{c}}$. The Atomic Cluster Expansion (ACE) framework provides a systematic construction of such models by expanding the site energies in terms of symmetric polynomials, organized by correlation order (or, equivalently, body order) as follows: For a given maximal correlation order $N_{\rm max}\in\mathbb{N}$, the ACE potential is formulated as
@@
Here, $Y_{\bm{\ell}}^{\bm{m}}$ denote the complex or real spherical harmonics, while $R_{\bm{n}\bm{l}}$ are radial basis functions with significant freedom of choice. Detailed implementation aspects, including the choice of the radial basis $R_{\bm{n}\bm{l}}$, can be found in [52] and Appendix -A.
@@
$\mathbf{B}:=\Big{\{}B_{{\bm{n}}{\bm{\ell}}q}\,\Big{|}\,$ $({\bm{n}},{\bm{\ell}})\in\mathbb{N}^{2N}\text{ (ordered)},\;\sum_{\alpha}\bm{\ell}_{\alpha}\text{ even},\;$
$\sum_{\alpha}m_{\alpha}=0,\;q=1,\ldots,n_{{\bm{n}}{\bm{\ell}}},\quad$
$\sum_{\alpha}(\bm{\ell}_{\alpha}+\bm{n}_{\alpha})\leq D_{\rm tot},\quad N\leq N_{\text{max}}\Big{\}}.$
@@
To estimate the coefficients $\bm{c}$, a training dataset of atomic configurations =$\mathfrak{R}=\{\mathbf{R}\}$ is required. For each configuration $\mathbf{R}$, target potential energy $\mathcal{E}_{\mathbf{R}}$, forces $\mathcal{F}_{\mathbf{R}}$ and virials $\mathcal{V}_{\mathbf{R}}$ are provided, usually from DFT calculations. The model parameters are then obtained by minimizing a least-squares loss function:
$L(\bm{c})$ $=\sum_{\mathbf{R}\in\mathfrak{R}}\Bigl{(}w_{E,\mathbf{R}}^{2}\left|E(\bm{c};\mathbf{R})-\mathcal{E}_{\mathbf{R}}\right|^{2}$
$\qquad\qquad+\;w_{F,\mathbf{R}}^{2}\left|F(\bm{c};\mathbf{R})-\mathcal{F}_{\mathbf{R}}\right|^{2}$
$\qquad\qquad+\;w_{V,\mathbf{R}}^{2}\left|V(\bm{c};\mathbf{R})-\mathcal{V}_{\mathbf{R}}\right|^{2}\Bigr{)}.$
@@
II-B Sparse recovery methods
Least-squares (LS) problems, such as (II-A), arise in numerous applications across signal processing, data science, and machine learning [32, 26]. In a standard LS problem, the objective is to solve for a vector $x\in\mathbb{R}^{n}$ that minimizes the residual $\|Ax-b\|_{2}^{2}$, where $A\in\mathbb{R}^{m\times n}$ is a known matrix, often termed a dictionary, and $b\in\mathbb{R}^{m}$ is the target vector. Sparse solutions, where most entries in $x$ are zero or near-zero, are particularly desirable in scenarios where only a small subset of predictors or signal components is relevant. Sparse solutions have important implications in model selection as they enable the identification of minimal predictor subsets that improve both interpretability and model generalization [44]. In signal processing, they facilitate signal reconstruction from limited measurements by assuming the signal is representable in a sparse basis [35]. Compressed sensing leverages sparsity to efficiently acquire and reconstruct signals from a reduced set of samples, thereby reducing data acquisition costs [11].
In optimization terms, the goal is to identify the smallest possible subset of features from the set represented by the columns of the matrix $A$. A sparse solution $x$ thus corresponds to selecting the minimal set of features that best approximates the target vector $b$. Formally, this sparse approximation problem, referred to as the Sparse Approximate Solution (SAS) problem, can be stated as follows: Given a matrix $A\in\mathbb{R}^{m\times n}$, where $A$ has $m$ rows and $n$ columns, a target vector $b\in\mathbb{R}^{m}$, and a tolerance $\epsilon>0$, find a vector $x$ that satisfies
@@ II-B Sparse recovery methods
and has the fewest possible nonzero entries among all vectors satisfying this condition. The sparse approximation selection (SAS) problem is known to be NP-hard [36]. However, practical alternatives exist, often classified into three main categories: convex relaxations, non-convex optimization techniques, and greedy algorithms. In what follows we will briefly review a few of these alternatives.
II-B1 Convex relaxations
Convex relaxations techniques provide practical alternatives for solving SAS and under specific conditions, can effectively approximate or even recover the exact sparsest solution [12]. A common approach is to formulate the problem using the $\ell_{1}$-norm objective and linear constraints (Basis Pursuit, BP [16]) or as a regularized least-squares problem (Basis Pursuit Denoising, BPDN [17]), or as the Least Absolute Shrinkage and Selection Operator (LASSO) [45]. The duals of these problems can be formulated as conventional linear and quadratic programs [49]. Consider the following primal and dual optimization problems:
@@ II-B1 Convex relaxations
subject $\text{subject to}\quad Ax+\lambda y=b,$
[…]
subject $\text{subject to}\quad-\mathbf{1}\leq A^{T}y\leq\mathbf{1},$
where $\lambda\geq 0$ is a scalar parameter and $\mathbf{1}$ is a vector of ones. When $\lambda=0$, this formulation is equivalent to the BP problem. For $\lambda>0$, the problem represents the BPDN problem. The primal and dual problems are duals of each other in the sense that there exists an optimal pair $(x,y)$ that simultaneously satisfy the Karush-Kuhn-Tucker (KKT) [31, 30] conditions for both problems. Various algorithms have been proposed for solving BP and BPDN. The BP-simplex algorithm [17], inspired by the simplex method in linear programming [51], begins by selecting an initial basis $A^{(0)}$ consisting of $m$ linearly independent columns of $A$. At each iteration, a swap is made to optimize the objective function by exchanging one term in the basis with one term outside the basis.
Friedlander and Saunders [27] introduced the parameterization (3) of the BPDN problem, based on a dual active-set approach suitable for large-scale problems, as it requires only matrix-vector products. Their algorithm iteratively identifies the active set of constraints, solves a least-squares subproblem on those constraints, and updates the solution estimate by moving along a descent direction. The algorithm continues until it reaches a stationary point where the objective gradient is a linear combination of the gradients of the active constraints, indicating optimality. We describe this algorithm in more detail in Appendix -B.
@@
[…]
by selecting the column of A most correlated with the current residual. After selecting a column, OMP updates the solution by solving a least-squares problem using the selected columns and computes a new residual orthogonal to the selected columns. OMP continues this process, adding one column at a time, until a stopping criterion is met, such as achieving a desired residual norm or reaching a predetermined number of iterations. The details of OMP will be discussed in Appendix -D.
The Homotopy method [37] is a technique for solving sparse least-squares problems by solving a sequence of LASSO problems with increasing regularization parameters $\tau_{k}$:
[…]
Homotopy balance the $\ell_{1}$ norm of the solution against the least-squares fit. As the regularization parameter decreases, the solution becomes sparser. This algorithm proceeds by identifying breakpoints, which are specific values of $\tau_{k}$ where the set of non-zero variables changes. The solution is updated systematically as $\tau_{k}$ passes through each breakpoint. We describe the Homotopy algorithm in more detail in Appendix -C.
II-B2 Empirical Bayes methods
@@
$b=Ax+\bm{\epsilon},$
where the error $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\lambda^{-1}\mathbf{I}),$ the likelihood function is
@@ A Gaussian prior is placed on the regression coefficients:
$p(x\mid b,A,\alpha,\lambda)=\mathcal{N}(x\mid\hat{x},\bm{\Sigma}),$
with:
$\bm{\Sigma}$ $=(\alpha\mathbf{I}+\lambda A^{\top}A)^{-1},$
$\hat{x}$ $=\lambda\bm{\Sigma}A^{\top}b.$
Hyperparameters $\alpha$ and $\lambda$ are typically optimized via evidence maximization.
@@
$p(x\mid\bm{\alpha})=\mathcal{N}(x\mid\mathbf{0},\mathbf{\Gamma}^{-1}),$
@@
In the context of ACE MLIPs, ARD has been used within the Hyperactive Learning (HAL) framework [50]. After fitting the model, posterior samples of the model parameters are drawn, and a committee of ACE models is formed to efficiently approximate the uncertainty in energy and force predictions. This uncertainty is then used in the HAL loop to identify configurations where the model is least certain. For further details on the derivations and optimization, refer to Appendix -E.
III Results
We present a number of benchmark results to test the proposed parameter estimation methods. In all benchmarks, a dataset was divided into training, validation, and test sets. The best checkpoint model was identified as the one achieving the lowest error on the validation set. For models of size $t$, the best checkpoint corresponds to the model with the best validation performance while having at most $t$ basis functions in the active set. These checkpoint models were then used to evaluate performance on the test set. In this section, we refer to the homotopy LASSO solver and the Orthogonal Matching Pursuit solver, introduced in Section II-B1, respectively, as ASP and OMP. We employ a Julia implementation ActiveSetPursuit.jl [47] of the BPdual, Homotopy, and OMP algorithms, based on [29].
An artifact of the BPDN model is its tendency to shrink coefficients toward zero as a result of the $\ell_{1}$-regularization. To de-bias [54, 48] the model, we applied a post-processing step using truncated singular value decomposition (TSVD) to refine the final coefficients. TSVD mitigates this issue by projecting the solution onto a lower-dimensional subspace, filtering out small singular values that contribute to numerical instability while preserving the dominant structures in the solution. This post-processing step thus improves both the accuracy and robustness of the final coefficient estimates. Further details on the TSVD procedure and its effects on the results can be found in Appendix -F.

### v2

@@ 1 Introduction
For large linear MLIPs models, achieving a balance between model complexity, interpretability, and generalizability remains a challenge [25, 26]. Sparse selection of features can in principle lead to minimal yet informative descriptions. Such models tend to enhance interpretability, scalability, and generalization to unseen configurations. Several previous studies have explored sparsification strategies for linear MLIPs. An early example of sparsification in this context appears in the MTP model [51], where a matching pursuit algorithm with crossover and local search was used to minimize an objective with an $\ell_{0}$ regularization term, to train a potential for tungsten. While sparsity was not the primary focus, this work demonstrated that a minimal basis could be identified without sacrificing accuracy. Seko et al. [49, 50] introduced the use of the least absolute shrinkage and selection operator (LASSO) [54] to automatically select basis functions from a large set of analytical radial descriptors. They demonstrated for elemental metals Na and Mg that a sparse potential energy surface can achieve high accuracy for energies and forces, while reducing the number of basis functions compared to neural network fits. Their focus was on showing that sparsity improves efficiency without sacrificing accuracy, using single-element systems and a fixed descriptor form.
Benoit et al. [7] studied the transferability of a linearized potential for Au–Fe nanoparticles, using LASSO regression with the Least Angle Regression scheme (LARS) [28], together referred to as LassoLARS. They applied this method to select descriptors from a large candidate pool, varying the penalty parameter to control model complexity. They found that adding more descriptors improved training-set accuracy but often reduced accuracy on unseen configurations, and suggested combining descriptor families to mitigate this effect. The aim of their study was to measure and interpret the complexity–transferability trade-off for a single challenging system, not to develop a general-purpose sparse MLIPs fitting methodology. In a related context, Goldman et al. [33] employed $\ell_{1}$-regularized regression (LASSO/LARS) to sparsify the linear ChIMES many-body model.
Automatic relevance determination (ARD) [40], an empirical Bayes method that can also be viewed as a sparsification technique, was used in the context of active learning of ACE models [60].
In the present work, we explore the use of two alternative sparse optimization methods, active set basis pursuit (ASP) [30], orthogonal matching pursuit (OMP) [30, 46], and compare them against ARD and various ”dense” least squares solvers. ASP and OMP generate entire ”paths” of MLIP models with increasing complexity in a single run, showing how accuracy and sparsity change as more basis functions are added. These active-set solvers avoid manual hyperparameter tuning by adjusting the regularization parameter automatically at each iteration so that exactly one additional basis function is selected. This approach gives full control over model complexity, removes the need for repeated cross-validation, and provides finer resolution in the accuracy–complexity trade-off. Our focus is on demonstrating the practical utility of sparse regression for linear MLIPs, and ACE in particular, especially analyzing how model quality varies under different sparse regression strategies.
Our computational tests on three benchmark problems demonstrate several advantages. The active-set sparse solvers improve generalization, yielding lower test errors than non-sparse methods, while requiring minimal user intervention and hyperparameter tuning. Moreover, the selected features show no clear or predictable pattern, emphasizing the benefit of data-driven feature selection over heuristic or pre-defined approaches.
2 Methods
2.1 Machine Learning Interatomic Potentials (MLIPs)
@@
A common approach to constructing interatomic potentials is to decompose the total energy $E$ of the system as a sum of the site energies: $E=\sum_{i}\epsilon_{i}.$ In linear models, each site energy $\epsilon_{i}$ is expressed as a linear combination of invariant basis functions that encode information about the local environment (e.g., interatomic distances, angles, etc.) within a cutoff radius $r_{\text{c}}$.
The Atomic Cluster Expansion (ACE) framework provides a systematic construction of such models by expanding the site energies in terms of symmetric polynomials, organized by correlation order (or, equivalently, body order) as follows: For a given maximal correlation order $N_{\rm max}\in\mathbb{N}$, the ACE potential is formulated as
@@
Here, $Y_{\bm{\ell}}^{\bm{m}}$ denote the complex or real spherical harmonics, while $R_{\bm{n}\bm{l}}$ are radial basis functions with significant freedom of choice. Detailed implementation aspects, including the choice of the radial basis $R_{\bm{n}\bm{l}}$, can be found in [62] and Appendix 6.1.
@@
$\mathbf{B}:=\Big\{B_{{\bm{n}}{\bm{\ell}}q}\,\Big|\,$ $({\bm{n}},{\bm{\ell}})\in\mathbb{N}^{2N}\text{ (ordered)},\;\sum_{\alpha}\bm{\ell}_{\alpha}\text{ even},\;\;\sum_{\alpha}m_{\alpha}=0,\;q=1,\ldots,n_{{\bm{n}}{\bm{\ell}}},\quad$
$\sum_{\alpha}(\bm{\ell}_{\alpha}+\bm{n}_{\alpha})\leq D_{\rm tot},\;N\leq N_{\text{max}}\Big\}.$
@@
To estimate the coefficients $\bm{c}$, a training dataset of atomic configurations $\mathfrak{R}=\{\mathbf{R}\}$ is required. For each configuration $\mathbf{R}$, target potential energy $\mathcal{E}_{\mathbf{R}}$, forces $\mathcal{F}_{\mathbf{R}}$ and virials $\mathcal{V}_{\mathbf{R}}$ are provided, usually from DFT calculations. The model parameters are then obtained by minimizing a least-squares loss function
$L(\bm{c})$ $=\sum_{\mathbf{R}\in\mathfrak{R}}\Bigl(w_{E,\mathbf{R}}^{2}\left|E(\bm{c};\mathbf{R})-\mathcal{E}_{\mathbf{R}}\right|^{2}+\;w_{F,\mathbf{R}}^{2}\left|F(\bm{c};\mathbf{R})-\mathcal{F}_{\mathbf{R}}\right|^{2}+\;w_{V,\mathbf{R}}^{2}\left|V(\bm{c};\mathbf{R})-\mathcal{V}_{\mathbf{R}}\right|^{2}\Bigr).$
@@
2.2 Sparse recovery methods
Least-squares (LS) problems, such as (2.1), arise in numerous applications across signal processing, data science, and machine learning [39, 29]. In a standard LS problem, the objective is to solve for a vector $x\in\mathbb{R}^{n}$ that minimizes the residual $\|Ax-b\|_{2}^{2}$, where $A\in\mathbb{R}^{m\times n}$ is a known matrix, often termed a dictionary, and $b\in\mathbb{R}^{m}$ is the target vector. Sparse solutions, where most entries in $x$ are zero or near-zero, are particularly desirable in scenarios where only a small subset of predictors or signal components is relevant. Sparse solutions have important implications in model selection as they enable the identification of minimal predictor subsets that improve both interpretability and model generalization [55]. In signal processing, they facilitate signal reconstruction from limited measurements by assuming the signal is representable in a sparse basis [42]. Compressed sensing leverages sparsity to efficiently acquire and reconstruct signals from a reduced set of samples, thereby reducing data acquisition costs [12].
In optimization terms, the goal is to identify the smallest possible subset of features from the set represented by the columns of the matrix $A$. A sparse solution $x$ thus corresponds to selecting the minimal set of features that best approximates the target vector $b$. The Sparse Approximate Solution (SAS) problem can be formulated as follows: Given a matrix $A\in\mathbb{R}^{m\times n}$, where $A$ has $m$ rows and $n$ columns, a target vector $b\in\mathbb{R}^{m}$, and a tolerance $\epsilon>0$, find a vector $x$ that satisfies
@@ 2.2 Sparse recovery methods
and has the fewest possible nonzero entries among all vectors satisfying this condition. The SAS problem is known to be NP-hard [44]. However, practical alternatives exist, often classified into three main categories: convex relaxations, non-convex optimization techniques, and greedy algorithms. In what follows we will briefly review a few of these alternatives.
2.2.1 Convex relaxations
Convex relaxations offer tractable alternatives to SAS and, under certain conditions, can approximate or even recover the exact sparsest solution [13]. Standard formulations include Basis Pursuit (BP) [17], Basis Pursuit Denoising (BPDN) [18], and LASSO [54], whose duals reduce to linear or quadratic programs [59]. Consider the following primal and dual optimization problems:
@@ 2.2.1 Convex relaxations
$\text{subject to}\quad Ax+\lambda y=b,$
Dual: $\underset{y}{\text{maximize}}\quad b^{T}y-\frac{1}{2}\lambda\|y\|_{2}^{2}$
$\text{subject to}\quad-\mathbf{1}\leq A^{T}y\leq\mathbf{1},$
where $\lambda\geq 0$ is the regularization parameter. Setting $\lambda=0$ yields BP, while $\lambda>0$ gives BPDN. Optimal pairs $(x,y)$ satisfy the KKT conditions [37, 34].
Algorithms for these problems include the BP-simplex method [18, 61], which iteratively swaps basis elements, and the Homotopy method [45], which solves a sequence of LASSO problems with decreasing regularization parameters $\tau_{k}$:
$\underset{x}{\text{minimize}}\;\|Ax-b\|_{2}^{2}\quad\text{subject to}\quad\|x\|_{1}\leq\tau_{k},$
updating the solution at breakpoints where the active set changes (see Appendix 6.3).
@@
$\underset{x}{\text{minimize}}\;\|Ax-b\|_{2}^{2}\quad\text{subject to}\quad\|x\|_{0}\leq k$
by selecting the column of A most correlated with the current residual. After selecting a column, OMP updates the solution by solving a least-squares problem using the selected columns and computes a new residual orthogonal to the selected columns. OMP continues this process, adding one column at a time, until a stopping criterion is met, such as achieving a desired residual norm or reaching a predetermined number of iterations. The details of OMP will be discussed in Appendix 6.4. Friedlander and Saunders [30] introduced the parameterization (3) of the BPDN problem, based on a dual active-set approach suitable for large-scale problems, as it requires only matrix-vector products. Their algorithm iteratively identifies the active set of constraints, solves a least-squares subproblem on those constraints, and updates the solution estimate by moving along a descent direction. The algorithm continues until it reaches a stationary point where the objective gradient is a linear combination of the gradients of the active constraints, indicating optimality. We describe this algorithm in more detail in Appendix 6.2.
Remark I. In the context of fitting MLIPs, the central goal is to identify the most influential basis functions. The path of solutions as the regularization parameter varies is crucial because it reveals the order in which basis functions enter the model and how their coefficients evolve. By tracing the full solution path, we can select a model that achieves predictive accuracy while retaining only the minimal set of meaningful basis functions.
We found that only a limited number of software packages reproduce the full solution path. In the Julia language, the main available implementations are Lasso.jl and LARS.jl, alongside our implementation of the ASP algorithm, originally developed in MATLAB and now available in Julia as ActiveSetPursuit.jl. Our numerical experiments revealed that LARS.jl and Lasso.jl lack robustness, efficiency, and practical usability, and they frequently fail to produce reliable results. This is demonstrated in Section 3.1.1.
2.2.2 Empirical Bayes methods
@@
$b=Ax+\boldsymbol{\epsilon},$
where the error $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\lambda^{-1}\mathbf{I}),$ the likelihood function is
@@ A Gaussian prior is placed on the regression coefficients:
$p(x\mid b,A,\alpha,\lambda)=\mathcal{N}(x\mid\hat{x},\boldsymbol{\Sigma}),$
with $\boldsymbol{\Sigma}=(\alpha\mathbf{I}+\lambda A^{\top}A)^{-1},$ and $\hat{x}=\lambda\boldsymbol{\Sigma}A^{\top}b.$ Hyperparameters $\alpha$ and $\lambda$ are typically optimized via evidence maximization.
@@
$p(x\mid\boldsymbol{\alpha})=\mathcal{N}(x\mid\mathbf{0},\mathbf{\Gamma}^{-1}),$
@@
In the context of ACE MLIPs, ARD has been used within the Hyperactive Learning (HAL) framework [60]. After fitting the model, posterior samples of the model parameters are drawn, and a committee of ACE models is formed to efficiently approximate the uncertainty in energy and force predictions. This uncertainty is then used in the HAL loop to identify configurations where the model is least certain. For further details on the derivations and optimization, refer to Appendix 6.5.
Remark II. Sparse regression, as considered in this work, refers to selecting a small subset of basis functions from a large candidate pool to build linear models. It should not be confused with sparse Gaussian process regression [38], where sparsity refers to data sparsity, achieved by using a limited set of inducing points to approximate the kernel and make training feasible on large datasets. It should also not be confused with tensor sketching [22] which is occasionally also referred to as a form of sparsification but is an entirely different form of dimension reduction.
3 Results
We present a number of benchmark results to test the proposed parameter estimation methods. In this section, we refer to the homotopy LASSO solver and the Orthogonal Matching Pursuit solver, introduced in Section 2.2.1, respectively, as ASP and OMP. We employ a Julia implementation ActiveSetPursuit.jl [57] of the BPdual, Homotopy, and OMP algorithms, based on [32].
An artifact of the ASP(BPDN) model is its tendency to shrink coefficients toward zero as a result of the $\ell_{1}$-regularization. To de-bias [64, 58] the model, we applied a post-processing step using truncated singular value decomposition (TSVD) to refine the final coefficients. TSVD mitigates this issue by projecting the solution onto a lower-dimensional subspace, filtering out small singular values that contribute to numerical instability while preserving the dominant structures in the solution. This post-processing step thus improves both the accuracy and robustness of the final coefficient estimates. Further details on the TSVD procedure and its effects on the results can be found in Appendix 6.6. Selecting the truncation parameter for the TSVD post-processing requires a validation set. The dataset splitting procedure is described in the following subsections.